OpenStudy (anonymous):

Helppp! If r is directly proportional to s^2 and s=(2/3) when r=(1/9). What is the positive value of s when r=(25/4)? A: 5/24 B: 5/16 C: 25/16 D: 5 E: 25

OpenStudy (anonymous):

its simple dude, use it as r1/r2 = (s1/s2)^2 here r1=1/9 s1=2/3 r2=25/4 find s2 Hope u understood

OpenStudy (anonymous):

i get it but why do i square (s1/s2) ?

OpenStudy (anonymous):

ohhh never mind i understand now! thank you very much!!!!!

OpenStudy (anonymous):

:D welcome